OpenStudy (henryblah):

If g is non-zero. How do I PROVE if (|g|)^0.5=g^m for some integer m, the only solutions are g=1 or -1. I know this is true just by looking at it but am struggling to write out a proof.

OpenStudy (amistre64):

perhaps assume it is false and show that such a notion is absurd

OpenStudy (amistre64):

or try some algebra? depends on what class this is in (|g|)^0.5=g^m

OpenStudy (amistre64):

\[ (|g|)^{0.5}=g^m\] \[ g^{0.5}=g^m\] \[ 1=g^{m-0.5}\]

OpenStudy (henryblah):

I'm trying to do a guided proof on why R* is not a cyclic group, this is one of the steps. Just not sure how rigorous I have to be.

OpenStudy (amistre64):

by an exponential property: g^0 = 1 for all g which means that m-.05 is never 0, so this cannot be true for all the other numbers that are not 1 .. but that is not a proof is it, its just saying that 1 is the only hopeful solution

OpenStudy (amistre64):

ah, group theory

OpenStudy (henryblah):

If 1 was the only answer, I would be fine with that answer there, but -1 works also, so it doesn't quite fit.

OpenStudy (amistre64):

the absolute value bars define a distance of 1, and distance is never negative

OpenStudy (amistre64):

bad wording .. the absolute value bars define a distance. and distance is never negative. |-1| = 1

OpenStudy (henryblah):

Ahh I think I see.

OpenStudy (henryblah):

Although that g on the right hand side...

OpenStudy (henryblah):

1=g^(m−0.5) implies that -1 can't be a solution does it not? because m-0.5 can't be any even number.

OpenStudy (amistre64):

we can algrebra this into: \[\huge 1=g^{\frac{2m-1}{2}}\] for any nonzero real number g we have g^0 = 1 by well known exponent properties. 2m-1 is an odd number, but 0 is an even number, so this excludes any real number .... leaving g=1 or -1 as test subjects

OpenStudy (amistre64):

not "any" real number ... but it narrows the dataset to -1 and 1

OpenStudy (henryblah):

Ok, I think i'll say that it leaves -1 and 1 to be checked, thought I somehow feel this isn't the neatest way to do it. Thanks for the help!

OpenStudy (amistre64):

lol, i never really have the "neatest" way to approach things ... thats the trouble when you have to make things up on the fly

OpenStudy (amistre64):

but it is something to work with :)

OpenStudy (amistre64):

\[(-1)^m=1\] for all even values of m